# Text-centric Alignment for Multi-Modality Learning

### v1

@@ 4.4.1. Text Transformation
Compared to the embedding-based methods SDEdit, analysis from Table 4 shows converting modality features into text enhances performance by approximately 2%, indicating less modality mismatch during training and inference compared to embedding representations. This improvement is consistent across most data modalities, except for tabular data, which sees about a 10% decline. This discrepancy is attributed to the fixed format of tabular text transformation, highlighting a significant style gap with more fluid, human-like writing, particularly impacting tabular data’s inference performance.
@@ 4.4.3. Text-Style Translation across Modality
According to Table 4, text-style translation bridges training and inference phase gaps, with about 7% improvements from 0.321 to 0.343. This enhancement is particularly notable when the gap in textual style remains consistent across phases, as seen in the image-to-table scenarios. Such consistency aids in more accurate mapping function determination by the model.
4.4.4. Reasoning Augmentation
Incorporating reasoning augmentation, as indicated in Table 4, may not necessarily yield significantly better performance compared to other components. However, we have observed that it contributes to a more stable performance across different scenarios. The variance value with augmentation is substantially lower than that without it.
@@ 5.1. Visualization for Distribution Alignment
In Section 4.3, we have validated the effectiveness of text transformation in TAMML through experimental performance. Further, we visualized 1,400 data points in these modalities with their position-aware embeddings using UMAP (McInnes et al., 2018) in Fig. 4. The left figure illustrates the original distributions of image and text embeddings, while the right figure displays the corresponding distributions after the summarization module in TAMML . We observe that the distribution boundaries between image and text modalities become less distinct, which indicates they are closer in the semantic space. To be more precise, TAMML significantly reduces the average instance Euclidean distance between image and text in the semantic space from $10.213$ to $0.411$.
5.2. Effects of the Foundation Models
Some might argue that the exceptional performance of TAMML could be attributed to the superior GPT-based foundation models. To investigate this, we replaced the different foundation model in our architecture with smaller open-source models like LLaMA for text and Blip2 for image captioning. We conducted ablation studies focusing on the performance of four various open-source LLMs and four image foundation models. Specifically, we showed that our approach maintains strong performance even with a small model size. Table 5 showcases the results averaged across twelve training-inference modality combinations (detailed shown in Appendix C.2, Tables 9 and  10). The results suggest that using smaller foundation models does not necessarily result in significantly inferior performance within TAMML .
Figure 4. The left and right pictures of the figure illustrate the visualizations of embeddings for image and text data, respectively, before and after our processes.
LLMs
Pet  —  Acc $\uparrow$ LLaMA-13b LLaMA-70b Mixtral-8x7b GPT-3.5
Average 0.302 0.317 0.320 0.321
@@ 5.2. Effects of the Foundation Models
Table 5. Foundation model comparison: Each number presented here is an average derived from detailed experiments outlined in Appendix C.2. In general, we can infer that the foundation model has only a limited impact on TAMML .
@@ 6. Conclusion and Future Directions
Appendix A Prompt Templates
User Prompt: task description The following is a dataset about Pet Adoption Prediction. We will predict the speed at which a pet is adopted based on the pet’s listing on the dataset. Input contains tabular modality. We previously converted this modality into text, and we tried to use the combined text to do classification tasks. modality description I would like you to rewrite all items in the input as text that represents image modality That is, transfer tabular text format ”The (column meaning) is (item)” into a format that describes the image, i.e., pretending that there is an image in front of you, and you go to describe it. Ensure that: 1. All information should naturally blend together from different modalities, just like the contents of the same paragraph, and the order should be switched and blended but not missing any information. 2. Please start with ”The image” Few-Shot Examples (Optional): Input: The type of pet is a Cat. The name of pet is Kali. The age of the pet when listed (in months) is 3… Answer: The image shows an adorable grey and black striped cat wearing a tiny yellow crown with…
Figure 5. Example prompt for text-style translation across modalities.
User Prompt: task description The following is a dataset about Pet Adoption Prediction. We will predict the speed at which a pet is adopted based on the pet’s listing on the dataset. The input contains multiple modalities, which are text and image modalities. We previously converted this modality into text, and we tried to use the combined text to do classification tasks. I would like you to rewrite all items in the input as natural text. modality description Ensure that: 1. If the original table contains entries under the ”secondary breed” and ”color” columns that are marked as ”Unknown,” the conversion into natural text will omit any mention of those particular attributes. 2. Avoid fabricating or inventing any content! 3. All information should naturally blend together from different modalities, just like the contents of the same paragraph, and the order should be switched and blended but not missing any information. 4. Please start with ”This pet is”. One-Shot Examples (Generated by LLM in the first step): Input: The type of pet is Cat. The name of pet is Kali. The age of pet when listed (in months) is 3… Answer: This pet is Kali, an adorable grey and black striped cat who carries …
Figure 6. Example prompt for modality summarization.
User Prompt: Background Information Airbnb revolutionized travel by connecting guests with unique accommodations worldwide. Its diverse offerings, from cozy homes to exotic locales, cater to individual preferences. Airbnb offers a range of accommodation options with varying prices, allowing users to find lodging that suits their budget and preferences. Prices are determined by hosts based on factors such as location, amenities, and property type. Task Description Imagine you are an assistant booking Airbnb, and you need to assist me since I want to predict a suitable price range for each Airbnb. I will provide you with some Airbnb examples, and you need to answer my question about the price of the given Airbnb information. For every question, explain as detailed as possible so I can evaluate the Airbnb value reasonably. Few-Shot Examples Following are examples of different Airbnb items and their price. Example 1: Input: The host’s response time to booking inquires… Answer: Price Range within \$$\$90$ to \$$\$110$ Example 2: … Inference Now, you are given information on Airbnb data. Sample Input: The detailed description of listing Airbnb is …. What price range do you think Airbnb falls within? Give me your answer and elaborate on your reason.
Figure 7. Example prompt for LLM reasoning augmentation.
Appendix B Experiment Detail Setup
B.1. Model Checkpoints
@@ B.2. Hyperparameters
SDEdit see folder in https://github.com/0Frett/LLM_Prediction_Baselines
DDRM see folder in https://github.com/0Frett/LLM_Prediction_Baselines
@@ B.4. Foundation Models
Appendix C Analysis and Discussion
C.1. In-context Modality Transfer Outperforms Zero-shot Learning Based Methods
@@
C.2. Effects of Foundation Model
To assess the impact of different foundational model capabilities in our architecture, we conducted ablation studies focusing on the performance of various LLMs and foundation models.
C.2.1. Large Language Model
In our experiments, LLMs primarily performed the bulk of text transformations. Therefore, the generation quality of these LLMs merits discussion. For this purpose, we compared the performance of LLaMA_13b, LLaMA_70b, Mixtral, and our baseline results in a simplified workflow focused solely on summarization tasks. The outcomes are detailed in Table 9. The data demonstrates that LLaMA_70b consistently surpasses LLaMA_13b, highlighting the influence of model size on summarization quality. However, when juxtaposed with Mixtral and our GPT-3.5 in various combinations, each model displayed unique strengths. Notably, LLaMA_70b showed superior performance in generalizing human-written texts. In contrast, Mixtral was more adept at processing texts related to images, while GPT-3.5 excelled in dealing with structured tabular data. These findings suggest that choosing the appropriate model based on the specific data modality can lead to more effective outcomes.
C.2.2. Image Caption Model
The quality of image captions significantly influences the outcomes of text transformation processes. In our simplified workflow, we focus on converting text with various image caption models. For this experiment, we selected three distinct image caption models: BLIP2 (Li et al., 2023), Kosmos2, and Flamingo. Each of them has undergone specialized pretraining for image captioning. As shown in Table 10, it is evident that the GPT-4 model maintains the highest stability across all mismatch combinations with at least a 30% accuracy rate. In contrast, other models drop to approximately 28% accuracy in tabular modality. However, in other combinations, the performance of these models is observed to be at similar levels. This indicates the feasibility of replacing image caption models in such frameworks.
Training Testing LLaMA-13b LLaMA-70b Mixtral-8x7b GPT-3.5
text+image tabular 0.315 0.316 0.336 0.321
text+tabular image 0.302 0.304 0.320 0.365
image+tabular text 0.300 0.308 0.295 0.295
text image+tabular 0.280 0.313 0.298 0.296
text image 0.318 0.320 0.311 0.298
text tabular 0.268 0.311 0.297 0.318
image text+tabular 0.322 0.300 0.341 0.289
image text 0.311 0.323 0.314 0.330
image tabular 0.316 0.332 0.334 0.307
tabular text+image 0.300 0.335 0.352 0.366
tabular text 0.292 0.308 0.308 0.306
tabular image 0.299 0.336 0.332 0.357
Average 0.302 0.317 0.320 0.321
Table 9. Compare different LLMs: The results show GPT-3.5 provides the most stable performance, while other models display unique strengths in different modality combinations, substantiating the feasibility of LLM replacement.
Training Testing Blip2 Kosmos2 Flamingo GPT4
image text+tabular 0.313 0.314 0.291 0.314
text 0.311 0.303 0.308 0.306
tabular 0.284 0.281 0.280 0.300
Average 0.303 0.299 0.293 0.307
Table 10. Compare different image caption models: The results show that GPT-4 provides the most stable performance, especially under tabular data, but under other combinations, all image caption models can achieve similar performance levels, substantiating the feasibility of image caption model replacement.
Appendix D Extra Experiment Results
@@
D.1. Q4: Text-based Solutions Versus Embedding-based Solutions When Training And Testing Modalities Are Identical
@@
D.2. Q5: How does TAMML compare to non-zero-shot methods?
@@
text image 0.288 0.350 0.195 0.338
text tabular 0.289 0.342 0.281 0.359
image text 0.270 0.324 0.276 0.306
image tabular 0.273 0.364 0.276 0.359
tabular text 0.289 0.326 0.195 0.306
tabular image 0.279 0.357 0.195 0.338
Table 12. The experiment results showed a condition with no train/test modality mismatch. Under this condition, TAMML shows no performance degradation and even performs better in several modality combinations. The regular method means the downstream model is trained on embedding representations. Note that this result differs from the one in Table 3 because the foundation models used for generating embedding and text representations are not the same.

### v2

@@ 4.4.1. Text Transformation
Compared to the embedding-based methods SDEdit, analysis from Table 4 shows converting modality features into text enhances performance by approximately 2%, indicating less modality mismatch during training and inference compared to embedding representations. This improvement is consistent across most data modalities, except for tabular data, which sees a decline of about 10%. This discrepancy is attributed to the fixed format of tabular text transformation, highlighting a significant style gap with more fluid, human-like writing, particularly impacting tabular data’s inference performance.
@@ 4.4.2. Modality Summarization
4.4.3. Reasoning Augmentation
Table 4 indicates that augmentation enhanced our average performance from 0.321 to 0.334. Additionally, we have observed that it contributes to a more stable performance across different scenarios. The variance value with augmentation is substantially lower than that without it.
@@ 4.4.4. Text-Style Translation across Modality
According to Table 4, text-style translation bridges training and inference phase gaps, with about 6% improvement from 0.334 to 0.355. This enhancement is particularly notable when the gap in textual style remains consistent across phases, as seen in the image-to-table scenarios. Such consistency aids in more accurate mapping function determination by the model.
@@ 5.1. Visualization for Distribution Alignment
In Section 4.3, we have validated the effectiveness of text transformation in TAMML through experimental performance. Furthermore, we visualized 1,400 data points in these modalities with their position-aware embeddings using UMAP (McInnes et al., 2018) in Figure. 5. The left figure illustrates the original distributions of image and text embeddings, while the right figure displays the corresponding distributions after the summarization module in TAMML . We observe that the distribution boundaries between image and text modalities become less distinct, which indicates they are closer in the semantic space. To be more precise, TAMML significantly reduces the average instance Euclidean distance between image and text in the semantic space from $10.213$ to $0.411$.
5.2. Effects of the Image Caption Models
Some might argue that the improvement in text transformation in TAMML could be attributed to the superior GPT-4 model. To investigate this, we replaced the different image caption models in our architecture with smaller open-source models. We conducted ablation studies focusing on the performance of four image foundation models. Specifically, we showed that our approach maintains strong performance even with smaller models. Table 5 showcases the results averaged across twelve training-inference modality combinations. The results suggest that using smaller image caption models does not necessarily result in significantly inferior performance with TAMML .
Figure 5. The left and right pictures illustrate the visualizations of embeddings for image and text data, respectively, before and after our processes.
@@ 5.2. Effects of the Image Caption Models
Table 5. Image caption model comparison: Each number presented here is an average derived from twelve modality combination experiments. In general, we can infer that the foundation model has only a limited impact on TAMML .
@@ 6. Conclusion and Future Directions
Appendix A Experiment Detail Setup
A.1. Model Checkpoints
@@ A.2. Hyperparameters
SDEdit batch_size=1, sample_step=3, noise_scale=150
DDRM batch_size=1, degredation_type=deno, noise=1.5
@@ A.4. Foundation Models
Appendix B Analysis and Discussion
B.1. In-context Modality Transfer Outperforms Zero-shot Learning Based Methods
@@
Appendix C Extra Experiment Results
@@
C.1. Q4: Text-based Solutions Versus Embedding-based Solutions When Training And Testing Modalities Are Identical
@@
C.2. Q5: How does TAMML compare to non-zero-shot methods?
@@
text image 0.288 0.374 0.195 0.338
text tabular 0.289 0.357 0.281 0.359
image text 0.270 0.319 0.276 0.306
image tabular 0.273 0.348 0.276 0.359
tabular text 0.289 0.364 0.195 0.306
tabular image 0.279 0.364 0.195 0.338
Table 10. The experiment results showed a condition with other non-zero-shot methods. Under this condition, TAMML shows no performance degradation and even performs better in several modality combinations in zero-shot.